Think about a monopolist, the market (inverse) demand function is: P = 30-2Q, his cost function is: C(Q) = 5+ Q2

  1. What is the monopolist’s optimal quantity and price? 
  2. What is the monopolist’s highest profit? 
1
Expert's answer
2021-12-14T09:54:47-0500

Solution:

1.). Profit maximizing quantity is where MR = MC

Derive MR:

TR = P "\times" Q

TR = (30 – 2Q) Q = 30Q – 2Q2

MR = "\frac{\partial TR} {\partial Q}" = 30 – 4Q

 

MC = "\frac{\partial TC} {\partial Q}" = 2Q

Set MR = MC

30 – 4Q = 2Q

30 = 2Q + 4Q

30 = 6Q

Q = 5

Optimal quantity = 5 units

Substitute in the demand function to derive optimal price:

P = 30 – 2Q

P = 30 – 2(5) = 30 – 10 = 20

P = 20

Optimal price = 20

 

2.). Monopolist highest profit = TR – TC

TR = P "\times" Q = 20 "\times" 5 = 100

TC = 5+ Q2 = 5 + 52 = 5 + 25 = 30

Monopolist highest profit = 100 – 30 = 70

Monopolist highest profit = 70